electric flux calculation
1
Expert's answer
2020-07-27T13:59:23-0400

For an uniform electric field


"\Phi=E\cdot S\cdot\cos \alpha"


For a non-uniform electric field


"\Phi=\int E\cdot \cos \alpha \cdot dS"